In Ryan Corporation, the first shift produced 5 1/2 times as many lightbulbs as the second shift. If the total lightbulbs produced was 16,250, how many lightbulbs were produced on shift #1?

Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)

x+5.5x=16250
6.5x=16250
x=2500
5.5x=13,750

Answer by richwmiller(17219)

f=5.5x
s=x
x+5.5x=16250
6.5x=16250
x=16250/6.5
solve for f
f=13750., s=2500., x=2500.